OpenStudy (anonymous):

Which equations are equivalent to the given equation? 49=-7k

OpenStudy (anonymous):

A.\[49\div(-7)=-7k \div(-7)\] B. \[49\times(-7)=-7k \div(-7)\] C. \[49\times(-\frac{ 1 }{ 7 })=-7k \times(-7)\] D. \[49\times(-\frac{ 1 }{ 7 })=-7k \times(-\frac{ 1 }{ 7 }\]

OpenStudy (anonymous):

\[k=-7\]

OpenStudy (anonymous):

It's either A or B, and I'm thinking it's A.

OpenStudy (anonymous):

I'm going with A.